A suspended simple pendulum of length l is making an angle θ with the vertical. On releasing, its velocity at lowest point will become?

Grade:12

2 Answers

Chandan
7 years ago
OC = OAcos theta = lcos theta h = BC=OB-OC= l (1-cos theta)K.E at B =P.A at A1/2mv^2=nutv^2=2gh=2gl (1-cos theta)v= root 2gl (1-cos theta)
Chandan
7 years ago
OC = OAcos theta = lcos theta h = BC=OB-OC= l (1-cos theta)K.E at B =P.A at A1/2mv^2=mghv^2=2gh=2gl (1-cos theta)v= root 2gl (1-cos theta)